OpenStudy (anonymous):

PLEASE HELP!!!! After spending (1/6) of my paycheck on bills, and (1/5) on rent, $350 is left. How much is my paycheck? Please also explain how you got the answer. Thanks! :)

jimthompson5910 (jim_thompson5910):

Let x = amount of paycheck Add the two fractions: 1/6+1/5 = 5/30 + 6/30 = 11/30 So you've spent 11/30 of your paycheck on bills and rent. This means that you've spent (11/30)*x dollars on bills and rent. If you spent 11/30 of your paycheck on bills and rent, and $350 is left, then this means that you spent x-350 dollars on bills and rent So equate the two (11/30)*x = x - 350 and now solve for x (11/30)*x = x - 350 30*(11/30)*x = 30x - 30*350 11x = 30x - 10500 11x - 30x = -10500 -19x = -10500 x = -10500/(-19) x = 552.63 So your paycheck is roughly 552.63 dollars

OpenStudy (anonymous):

Wow! Thank you so much! You're definitely good at explaining problems. Keep up the great work!

jimthompson5910 (jim_thompson5910):

why thank you, glad to be of help

OpenStudy (anonymous):

:)